# Paul-Erik Haacker, Remco I. Leine, Renu Chaudhary, Kai Diethelm, André Schmidt, Safoura Hashemishahraki

### v1

@@ Abstract
This work has been submitted to the Journal of Nonlinear Dynamics (Springer) and is currently under review.
@@ 1 Introduction
Recently, interest in fractional-order differential equations has grown in various fields. A number of dynamical processes with memory have proven to be better described using fractional-order differential equations, where integer-order models fail to provide low-degree-of-freedom representations. Among others, applications include viscoelastic materials, diffusion processes and control theory [11]. Such – possibly nonlinear – systems admit periodic solutions when periodically forced or in the presence of self-excitation. Common formulations of Caputo differential equations cannot have exactly periodic solutions [14], which is closely related to the fact that Caputo derivatives and Riemann-Liouville integrals of periodic functions are only asymptotically periodic [20]. Studies [12, 13] concerned with finding steady-state solutions of differential equations involving Liouville-Weyl-type fractional-order derivatives have successfully applied harmonic balance methods. A linearization of the system dynamics about a periodic solution, describing the dynamics of small perturbations, yields a LTP-FODE. This motivates the study of the stability of such systems.
@@
Lyapunov-stability of equilibria and $\Omega$-limit sets is a central concept in the study of dynamical systems and has also been studied for fractional order systems. Classically, stability of linear time-invariant fractional-order differential equations can be studied through Mittag-Leffler functions, see [2] and the references cited therein. As fractional-order systems typically admit algebraic decay [25], the notion of Mittag-Leffler stability was introduced analogously to exponential stability [28]. Furthermore, as fractional-order derivatives are nonlocal, FODEs are similar to functional differential equations [22]. This similarity has been exploited by [6, 16] in mechanical systems with fractional-order damping, treating the system as a functional differential equation with infinite delay and finding Lyapunov-functionals through the infinite-state representation (also called diffusive representation or kernel compression), which is often used for efficient numerical approximations of fractional-order integrals [10]. LTI-FODEs perturbed by small nonlinearities were studied in [24]. That work also analyzes linear time-varying FODEs, but only under the assumption of sufficiently small time-dependent terms. To the best of our knowledge, no stronger results on stability of solutions of linear time-variant fractional-order Caputo differential equations are available.
@@
In classical integer-order models (i.e. first-order ordinary differential equations), stability of periodic solutions may be analyzed via the eigenvalues of the infinite-dimensional Hill matrix which correspond to the Floquet exponents, see [4] and the references cited therein. These problems are understood quite deeply, with Floquet theory as their foundation. This theory may be applied in bifurcation analysis of periodic motion of mechanical systems, which is self-exited [31] or forced in the form of frequency response curves [30]. For LTP-FODEs, no similar theory has been established. When employing an approach analogous to that used for first-order systems, many difficulties arise for fractional-order systems [18]. To reach a Floquet theory, results on fundamental solution operators for LTP systems are lacking. The more general case of linear time-varying systems was considered by [17], developing a fundamental solution matrix. Defining a monodromy operator and analyzing its spectrum, as done for functional differential equations [22], remains an open challenge in the study of LTP-FODEs. Previous studies performed transition curve analysis on the fractional-order Mathieu equation [12, 21], as well as on more general forms  [13], also including first-order derivatives and time-delays. The latter work employs fractional-order derivatives with infinite memory, i.e. Liouville-Weyl-type derivative, which is of particular relevance to our study. However, the analysis of transition curves in the parameter space only gives information on stability boundaries and thus fails to completely characterize stability over a parameter space, nor does it give information about the type of bifurcation which may originate at the stability boundary.
@@
The literature review shows, that the classical Caputo-type differential equations do not admit exactly periodic solutions, but do admit them as limit sets. In order to study the stability properties of such limit sets, one must first formulate an extension of the class of Caputo differential equations to also include periodic solutions. The Liouville-Weyl derivative has these properties and a rigorous formulation of an initial value problem along with the set of initial conditions has to be developed. This then opens paths to extend classical theories on periodic solutions to fractional-order systems. For example, fundamental tools such as the Poincaré–Bendixson Theorem [23], which provides conditions for the existence of periodic solutions, need to be developed in the fractional-order case. Further, when given a periodic solution, possibly obtained using the harmonic balance method [12, 13, 21], linearly perturbed dynamics can be studied to assess stability in the hyperbolic case. Therefore, a Floquet-type theory on LTP-FODEs needs to be developed and the existence of Floquet exponents needs to be studied. For application of this theory to numerical stability analysis, methods like the Hill method should be extended to the fractional-order case.
@@
The main merit of this paper is the insight that exponentially decaying Floquet solutions (i.e. the ones related to asymptotically stable equilibria or stable manifolds of saddle equilibria) do not exist for LTP-FODEs, but exponentially growing Floquet solutions may be observed and assessed via a Hill-type approach. Parts of the research of this work were presented in a preliminary form at the International Conference on Nonlinear Dynamics and Applications 2025 [29], namely the fractional Hill problem was presented as well as results on linear time-invariant FODEs. The novel contributions of the present work are the formulation of the system class in Section 3, the analysis of decaying solutions in Section 4 as well as the proof of Theorem 5.1 and the following Propositions.
@@ 2.1 Fractional Calculus
We briefly recall some concepts of fractional calculus and refer the reader to the books [2, 26] for further details. Several definitions of fractional derivatives exist, each suited to different mathematical and physical contexts. Given a function $f\in L^{1}[a,b]$, the Riemann–Liouville (RL) fractional integral of order $\alpha>0$ is defined as
@@
where $\Gamma(\cdot)$ is the Gamma function. The corresponding RL fractional derivative is given by
$({}^{\mathrm{RL}}\operatorname{D}_{a}^{\alpha}f)(t)=\frac{d^{n}}{dt^{n}}\left(I_{a+}^{n-\alpha}f\right)(t),$ (2)
with $n=\lceil\alpha\rceil$. The RL derivative requires fractional-order initial conditions, which limits its practical applicability in initial value problems.
To overcome this, the Caputo derivative is often used in applied settings [6, 12, 13]. For $\alpha\in(n-1,n)$ with some $n\in\mathbb{N}$ and a function $f\in A^{n}[a,b]$, it is defined as
$\left({}^{\mathrm{C}}\operatorname{D}_{a}^{\alpha}f\right)(t)=\frac{1}{\Gamma(n-\alpha)}\int_{a}^{t}(t-\tau)^{n-\alpha-1}f^{(n)}(\tau)\,d\tau,$ (3)
@@
$D^{\alpha}x(t):=\left(D^{\alpha}x_{1}(t),D^{\alpha}x_{2}(t),\dots,D^{\alpha}x_{n}(t)\right)^{\top},$ (4)
@@
$E_{\alpha,\beta}(Z)=\sum_{k=0}^{\infty}\frac{Z^{k}}{\Gamma(\alpha k+\beta)},\quad E_{\alpha}(Z)=E_{\alpha,1}(Z).$ (5)
In the case $\alpha=1$, the Mittag-Leffler function reduces to the exponential function, i.e., $E_{1}(z)=e^{z}$.
@@
$({}^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}f)(t)=\frac{1}{\Gamma(n-\alpha)}\int_{-\infty}^{t}(t-\tau)^{n-\alpha-1}f^{(n)}(\tau)\,d\tau,$ (6)
with $n=\lceil\alpha\rceil$. Here, we use the Caputo-type definition $^{\mathrm{C}}\operatorname{D}_{a}^{\alpha}\to{}^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}$ for $a\to$-$\infty$. One may show for sufficient assumptions on integrability and differentiability of $f$, that it holds that $^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}f={}^{\mathrm{RL}}\operatorname{D}_{-\infty}^{\alpha}f$. In this work, we will consider fractional orders $\alpha\in(0,1)$.
@@ 2.2 The Hill Method for First-Order Systems
Hill’s method is a frequency-based approach to assess the stability of periodic solutions [19]. It has been applied to first order ordinary differential equations as well as delay differential equations [27]. Here, we summarize the approach for ordinary differential equations, which we develop for FODEs in Section 5. For time-periodic first-order systems
@@
where $\tilde{f}(t,\cdot)=\tilde{f}(t+T,\cdot)$ for all $t$ and a $T>0$, there may exist a known periodic solution $\xi_{p}(t)=\xi_{p}(t+T)$ and desire to assess its stability properties. To do so, one can introduce the perturbation $\delta(t)=\xi(t)-\xi_{p}(t)$ and find its linearized dynamics to be
@@
where $\mu$ is called Floquet exponent and $\tilde{p}$ is $T$-periodic, the Floquet exponents are eigenvalues of the infinite-dimensional Hill-matrix
@@
Herein, $\tilde{J}_{k}$ are the Fourier coefficients of $\tilde{J}(t)=\sum_{k=-\infty}^{+\infty}\tilde{J}_{k}\exp(ik\omega t)$, $\omega=2\pi/T$ and $I$ is the identity matrix of appropriate dimension. After truncation to finite dimension $N$, one may either compute all eigenvalues of the truncated Hill-matrix $\mathbf{H}_{N}$ and employ a sorting technique to extract Floquet exponents or perform the recently introduced Koopman-Hill projection [4], avoiding sorting and directly approximating Floquet-multipliers.
@@ 3 Liouville-Weyl Differential Equations: A Class of Systems with Periodic Solutions
to form an initial value problem (11). Note that the solution $x:\mathbb{R}\to\mathbb{R}^{n}$ of (11) is required to satisfy the differential equation (11a) only for times $t\geq t_{0}$. The initialization function $x_{0}$ is assumed to be a member of function space $X(-\infty,t_{0}]$ given in Definition 3.1 below. Therein, the local boundedness of $x_{0}^{\prime}$ at $t_{0}$ is indeed necessary, as we will see in Example 3.6 (iv) below.
@@
$X(-\infty,t_{0}]:=$ $\Bigl\{x_{0}\in C(-\infty,t_{0}]\cap L^{\infty}(-\infty,t_{0}]~|~x_{0}^{\prime}\text{ exists a.e. and }\exists\,\eta>0:x_{0}^{\prime}\in L^{\infty}[t_{0}-\eta,t_{0}]\Bigr\}$ (12)
where $C(a,b]$ and $L^{\infty}(a,b]$ denotes the function spaces of continuous functions and functions with a finite supremum-norm, respectively on the interval $(a,b]$.
In order to prove existence and uniqueness of solutions of problem (11), we will split the Liouville-Weyl derivative
@@
into the Caputo derivative of lower bound $t_{0}$ and a forcing term $\mathcal{F}x_{0}$, which is given in Definition 3.2 below. The properties of the forcing term $\mathcal{F}x_{0}$ along with classical results on Caputo fractional differential equations allows to establish existence and uniqueness of solutions of problem (11). We formalize this after Lemma 3.4.
@@
to which we will refer to as the forcing term of the initial condition.
The forcing term $\mathcal{F}x_{0}(t)$ of the initial condition is continuous and has the following decay property, which we will use in the analysis of a linear time-invariant version of (11a) in Section 4. The decay bound may be regarded as a fading-memory property, where the differential equation ”forgets” events of the distant past.
@@
For an initial condition $x_{0}\in X(-\infty,t_{0}]$, its forcing term $\mathcal{F}x_{0}$ is continuous on $t_{0},\infty)$ and may be bounded
@@
The proof of Lemma 3.3 is given in the Appendix A. As motivated in equation (13), one may split the Liouville-Weyl derivative into Caputo derivative and forcing term. Thus, any solution of the Liouville-Weyl initial value problem (11) always satisfies a classical Caputo differential equation. We formalize this in the following Lemma.
Lemma 3.4.
A function $x\in C[t_{0},t_{0}+h)$ with $h>0$ is a solution of (11) if and only if it satisfies the equation
${}^{\mathrm{C}}\operatorname{D}_{t_{0}}^{\alpha}x(t)$ $=f(t,x(t))-\mathcal{F}x_{0}(t)=:f_{0}(t,x(t)),\quad t\geq t_{0},$ (17)
with the initial condition $x(t_{0})=x_{0}(t_{0})$.
Proof 3.5.
This is a direct consequence of equation (13).
The newly introduced $f_{0}$ generally carries an explicit time-dependence introduced by the forcing $\mathcal{F}x_{0}$, even when $f$ is not explicitly time-dependent. Existence results for initial value problems of the form (17) are available under suitable conditions on $f_{0}$, as discussed in [2]. These conditions are satisfied in the present case. In particular, since $f$ is assumed Lipschitz in its second argument, so will be $f_{0}$. As $\mathcal{F}x_{0}$ and $f$ are continuous, so is $f_{0}$.
@@
    The function $x_{0}(t)=\exp(-t)$ does not belong to $X(-\infty,0]$, as its forcing term does not exist for any $t\geq 0$. Indeed, it holds that
@@
    The function $x_{0}$ is continuous and bounded everywhere and its derivative exists everywhere except at t=1t=-1 and is bounded in the neighborhood of 0.

### v2

@@ Abstract
  • $\cdot$
    For a tuple $x=(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$, $x||_{2}=\sqrt{\sum_{k=1}^{n}x_{k}^{2}}$ denotes the 2-norm.
@@
  • $\cdot$
    $AC[a,b]$ denotes the space of functions $f:[a,b]\to\mathbb{R}$ which are absolutely continuous on $[a,b]$. We say that a function $f$ is locally absolutely continuous on $(-\infty,0]$, if it is absolutely continuous on every finite subinterval. We denote this $f\in AC_{\mathrm{loc}}(-\infty,0]$.
@@ 1 Introduction
Recently, interest in fractional-order differential equations has grown in various fields. A number of dynamical processes with memory have proven to be better described using fractional-order differential equations, where integer-order models fail to provide low-degree-of-freedom representations. Among others, applications include viscoelastic materials, diffusion processes and control theory [Das2011]. Such – possibly nonlinear – systems admit periodic solutions when periodically forced or in the presence of self-excitation. Common formulations of Caputo differential equations cannot have exactly periodic solutions [Tavazoei2009], which is closely related to the fact that Caputo derivatives and Riemann-Liouville integrals of periodic functions are only asymptotically periodic [Area2016]. Studies [Rand2010, Daibiri2015] concerned with finding steady-state solutions of differential equations involving Liouville-Weyl-type fractional-order derivatives have successfully applied harmonic balance methods. A linearization of the system dynamics about a periodic solution, describing the dynamics of small perturbations, yields a LTP-FODE. This motivates the study of the stability of such systems.
@@
Lyapunov-stability of equilibria and $\Omega$-limit sets is a central concept in the study of dynamical systems and has also been studied for fractional order systems. Classically, stability of linear time-invariant fractional-order differential equations can be studied through Mittag-Leffler functions, see [Diethelm2010] and the references cited therein. As fractional-order systems typically admit algebraic decay [CongEtAl2020], the notion of Mittag-Leffler stability was introduced analogously to exponential stability [LiEtAl2009]. Furthermore, as fractional-order derivatives are nonlocal, FODEs are similar to functional differential equations [Hale1977]. This similarity has been exploited by [Hinze2020, Hinze2020a] in mechanical systems with fractional-order damping, treating the system as a functional differential equation with infinite delay and finding Lyapunov-functionals through the infinite-state representation (also called diffusive representation or kernel compression), which is often used for efficient numerical approximations of fractional-order integrals [Chaudhary2024]. LTI-FODEs perturbed by small nonlinearities were studied in [CongEtAl2018]. That work also analyzes linear time-varying FODEs, but only under the assumption of sufficiently small time-dependent terms. To the best of our knowledge, no stronger results on stability of solutions of linear time-variant fractional-order Caputo differential equations are available.
In classical integer-order models (i.e. first-order ordinary differential equations), stability of periodic solutions may be analyzed via the eigenvalues of the infinite-dimensional Hill matrix which correspond to the Floquet exponents, see [Bayer2023a] and the references cited therein. These problems are understood quite deeply, with Floquet theory as their foundation. This theory may be applied in bifurcation analysis of periodic motion of mechanical systems, which is self-excited [Horvath2022] or externally forced [Youssef2025]. For LTP-FODEs, no similar theory has been established, although several attempts for the development of a Floquet theory for fractional-order systems have been undertaken [Rezazadeh2016, Iomin2023]. In particular, the work [Rezazadeh2016] assumes that the Mittag-Leffler function approximately satisfies a multiplicative identity, despite the fact that the associated error may become arbitrarily large. Further, in [Rezazadeh2016] a theory entirely based on approximations is developed, which only holds exactly in the first-order case ($\alpha=1$). However, parameter identification schemes applied to experimental data typically yield fractional orders between $0.1$ and $0.7$ [Pawlak2021, HinzeExperiment, Kim2009, Stark2021]. The work [Iomin2023] considers a Schrödinger LTP-FODE of Caputo-type. Therein, a solution form expressed by a series of Mittag-Leffler functions is considered; however, an analysis of asymptotic properties of the solution form as well as a Hill-type method to access characteristic multipliers is missing. Furthermore, the proposed form of solutions of [Iomin2023] uses a basis of Mittag–Leffler functions whereas the time-periodic system matrix (the Hamiltonian in the terminology of [Iomin2023]) is represented in harmonic form. The use of two different series impedes a Hill-type approach, since the Mittag-Leffler function and the exponential function do not share a multiplicative identity.
@@
To reach a Floquet theory, results on fundamental solution operators for LTP systems are lacking. The more general case of linear time-varying systems was considered by [Gomoyunov2020], developing a fundamental solution matrix. Defining a monodromy operator and analyzing its spectrum, as done for functional differential equations [Hale1977], remains an open challenge in the study of LTP-FODEs. Previous studies performed transition curve analysis on the fractional-order Mathieu equation [Rand2010, Shahroudi2024], as well as on more general forms  [Daibiri2015], also including first-order derivatives and time-delays. The latter work employs fractional-order derivatives with infinite memory, i.e. Liouville-Weyl-type derivatives, which is of particular relevance to our study. However, the analysis of transition curves in the parameter space only gives information on stability boundaries and thus fails to completely characterize stability over a parameter space, nor does it give information about the type of bifurcation which may originate at the stability boundary.
@@
The literature review shows, that the classical Caputo-type differential equations do not admit exactly periodic solutions, but do admit them as limit sets. In order to study the stability properties of such limit sets, one must first formulate an extension of the class of Caputo differential equations to also include periodic solutions. The Liouville-Weyl derivative has these properties and a rigorous formulation of an initial value problem along with the set of initial conditions has to be developed. This then opens paths to extend classical theories on periodic solutions to fractional-order systems. For example, fundamental tools such as the Poincaré–Bendixson Theorem [HirschEtAl2013], which provides conditions for the existence of periodic solutions, need to be developed in the fractional-order case. Further, when given a periodic solution, possibly obtained using the harmonic balance method [Rand2010, Daibiri2015, Shahroudi2024], linearly perturbed dynamics can be studied to assess stability in the hyperbolic case. Therefore, a Floquet-type theory on LTP-FODEs needs to be developed and the existence of Floquet exponents needs to be studied. For application of this theory to numerical stability analysis, methods like the Hill method should be extended to the fractional-order case.
@@
The main merit of this paper is the insight that exponentially decaying Floquet solutions (i.e. the ones related to asymptotically stable equilibria or stable manifolds of saddle equilibria) do not exist for LTP-FODEs, but exponentially growing Floquet solutions may be observed and assessed via a Hill-type approach. Parts of the research of this work were presented in a preliminary form in reference [Haacker2025], namely the fractional Hill problem was presented as well as results on linear time-invariant FODEs. The novel contributions of the present work are the formulation of the system class in Section 3, the analysis of decaying solutions in Section 4 as well as the proof of Theorem 5.1 and the following Propositions.
@@ 2.1 Fractional Calculus
We briefly recall some concepts of fractional calculus and refer the reader to the books [Diethelm2010, Kilbas2006] for further details. Several definitions of fractional derivatives exist, each suited to different mathematical and physical contexts. Given a function $f\in L^{1}[a,b]$, the Riemann–Liouville (RL) fractional integral of order $\alpha>0$ is defined as
@@
where $\Gamma(\cdot)$ is the Gamma function. In this work, we consider derivatives of fractional orders $\alpha\in(0,1)$. The RL fractional derivative of order $\alpha\in(0,1)$ is given by
$\left({}^{\mathrm{RL}}\operatorname{D}_{a}^{\alpha}f\right)(t)=\left(\operatorname{D}^{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}I_{a}^{1-\alpha}}f\right)(t).$ (2)
The RL derivative requires fractional-order initial conditions, which limits its practical applicability in initial value problems.
To overcome this limitation of the RL derivative, the Caputo derivative is often used in applied settings [Hinze2020, Rand2010, Daibiri2015]. For $\alpha\in(0,1)$ and a function $f\in AC[a,b]$, it is defined as
$\left({}^{\mathrm{C}}\operatorname{D}_{a}^{\alpha}f\right)(t)=\left({I_{a}^{1-\alpha}}\operatorname{D}^{1}f\right)(t)=\frac{1}{\Gamma(1-\alpha)}\int_{a}^{t}(t-\tau)^{-\alpha}f^{\prime}(\tau)\,d\tau,$ (3)
@@
$\operatorname{D}^{\alpha}x(t):=\left(\operatorname{D}^{\alpha}x_{1}(t),\operatorname{D}^{\alpha}x_{2}(t),\dots,\operatorname{D}^{\alpha}x_{n}(t)\right)^{\top},$ (4)
@@
A different formulation (see [Diethelm2010] and the references therein) of the Caputo derivative reads
$\left(\operatorname{D}_{\ast a}^{\alpha}f\right)(t)=\left({}^{\mathrm{RL}}\operatorname{D}_{a}^{\alpha}[f-f(0)]\right)(t),$ (5)
for the case of $\alpha\in(0,1)$. If $f\in AC[a,b]$, then the two formulations coincide $\operatorname{D}_{\ast a}^{\alpha}f={}^{\mathrm{C}}\operatorname{D}_{a}^{\alpha}f$ almost everywhere [Diethelm2010].
@@
$E_{\alpha,\beta}(Z)=\sum_{k=0}^{\infty}\frac{Z^{k}}{\Gamma(\alpha k+\beta)},\quad E_{\alpha}(Z)=E_{\alpha,1}(Z),$ (6)
respectively. In the case $\alpha=\beta=1$, the Mittag-Leffler function reduces to the exponential function, i.e., $E_{1}(z)=e^{z}$.
@@
$({}^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}f)(t)=\frac{1}{\Gamma(1-\alpha)}\int_{-\infty}^{t}(t-\tau)^{-\alpha}f^{\prime}(\tau)\,d\tau.$ (7)
Here, we use the Caputo-type definition $^{\mathrm{C}}\operatorname{D}_{a}^{\alpha}\to{}^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}$ for $a\to$-$\infty$. One may show for sufficient assumptions on integrability and differentiability of $f$, that it holds that $^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}f={}^{\mathrm{RL}}\operatorname{D}_{-\infty}^{\alpha}f$.
@@ 2.2 The Hill Method for First-Order Systems
Hill’s method is a frequency-based approach to assess the stability of periodic solutions [Zhou2002]. It has been applied to first order ordinary differential equations as well as delay differential equations [Stepan2011]. Here, we summarize the approach for ordinary differential equations, which we develop for FODEs in Section 5. For time-periodic first-order systems
@@
where $\tilde{f}(t,\cdot)=\tilde{f}(t+T,\cdot)$ for all $t$ with some $T>0$, there may exist a known periodic solution $\xi_{p}(t)=\xi_{p}(t+T)$ and the desire to assess its stability properties. To do so, one can introduce the perturbation $\delta(t)=\xi(t)-\xi_{p}(t)$ and find its linearized dynamics to be
@@
where $\mu$ is called the Floquet exponent and $\tilde{p}$ is $T$-periodic, the Floquet exponents are eigenvalues of the infinite-dimensional Hill-matrix
@@
Herein, $\tilde{J}_{k}$ are the Fourier coefficients of $\tilde{J}(t)=\sum_{k=-\infty}^{+\infty}\tilde{J}_{k}\exp(ik\omega t)$, $\omega=2\pi/T$ and $I$ is the identity matrix of appropriate dimension. After truncation to finite dimension $N$, one may either compute all eigenvalues of the truncated Hill-matrix $\mathbf{H}_{N}$ and employ a sorting technique to extract Floquet exponents or perform the recently introduced Koopman-Hill projection [Bayer2023a], avoiding sorting and directly approximating Floquet-multipliers.
@@ 3 Liouville-Weyl Differential Equations: A Class of Systems with Periodic Solutions
to form an initial value problem (12). For some $t_{1}>t_{0}$, we call a function $x:(-\infty,t_{1})\to\mathbb{R}^{n}$ a strong solution of the initial value problem (12) if it is continuous, coincides with the initial condition $x_{0}$ on $(-\infty,t_{0}]$ and satisfies the differential equation (12a) for $t_{0}<t<t_{1}$. The initialization function $x_{0}$ is assumed to be an element of the function space $X(-\infty,t_{0}]$ given in Definition 3.1 below. Therein, the local boundedness of $x_{0}^{\prime}$ at $t_{0}$ is indeed necessary, as we will see in Example 3.7 (iv) below.
@@
$X(-\infty,t_{0}]:=$ $\Bigl\{x_{0}\in{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}AC_{\mathrm{loc}}(-\infty,t_{0}]}\cap L^{\infty}(-\infty,t_{0}]~|~\exists\,\eta>0:x_{0}^{\prime}\in L^{\infty}[t_{0}-\eta,t_{0}]\Bigr\}$ (13)
where $AC_{\mathrm{loc}}(a,b]$ and $L^{\infty}(a,b]$ denote the function spaces of locally absolutely continuous functions and functions with a finite supremum-norm, respectively, on the interval $(a,b]$.
In order to assess the existence and uniqueness of solutions of problem (12), we will split the Liouville-Weyl derivative
@@
into the Caputo derivative of lower bound $t_{0}$ and a forcing term $\mathcal{F}x_{0}$, which is given in Definition 3.2 below. The properties of the forcing term $\mathcal{F}x_{0}$ along with classical results on Caputo fractional differential equations allows to establish existence and uniqueness of solutions of problem (12). We formalize this in Proposition 3.5.
@@
to which we will refer as the forcing term of the initial condition.
The forcing term $\mathcal{F}x_{0}(t)$ of the initial condition is continuous and has the following decay property, which we will use in the analysis of a linear time-invariant version of (12a) in Section 4. The decay bound may be regarded as a fading-memory property, where the differential equation“forgets” events of the distant past.
@@
For an initial condition $x_{0}\in X(-\infty,t_{0}]$, its forcing term $\mathcal{F}x_{0}$ is continuous on $t_{0},\infty)$ and may be bounded according to
@@
The proof of Lemma 3.3 is given in Appendix A. As motivated in equation (14), one may split the Liouville-Weyl derivative into Caputo derivative and forcing term. Thus, a strong solution of the Liouville-Weyl initial value problem (12) always satisfies a Caputo initial value problem
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{}^{\mathrm{C}}\operatorname{D}_{t_{0}}^{\alpha}x(t)$ $=\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f(t,x(t))-\mathcal{F}x_{0}(t)=:f_{0}(t,x(t)),\quad t\geq t_{0},\quad x(t_{0})=x_{0}(t_{0}).$ (18)
Typically, the $\operatorname{D}_{\ast 0}^{\alpha}$-type differential equations are considered instead of $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type formulations, since the latter shows problems on the space $C[0,h]$, but the two formulations coincide on $AC[0,h]$, see [Webb2025, Section 2.3]. We formalize this for our case in the Proposition 3.5 below. Consider the following example of $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type problems on $C[0,h]$.
Example 3.4.
Let $x:[0,1]\to\mathbb{R}$ be the Cantor ternary function [GO2003, pp. 96ff.]. This function is continuous, monotonically increasing, differentiable almost everywhere, and satisfies $x^{\prime}(t)=0$ a. e. Thus, $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x(t)=0$ for $t\in[0,1]$ and $x$ is a strong solution of the initial value problem $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x(t)=0$, $x(0)=0$. By a strong solution we mean a continuous function satisfying the differential equation and initial condition.
To exclude this odd behavior, we may instead study the $\operatorname{D}_{\ast 0}^{\alpha}$-type initial value problem
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\operatorname{D}_{\ast t_{0}}^{\alpha}x(t)$ $=\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f(t,x(t))-\mathcal{F}x_{0}(t)=:f_{0}(t,x(t)),\quad t\geq t_{0},\quad x(t_{0})=x_{0}(t_{0}).$ (19)
Since $f$ is continuous in its first variable and locally Lipschitz in its second variable, the same properties hold for $f_{0}$, because Lemma 3.3 guarantees that the forcing term $\mathcal{F}x_{0}$ is continuous. We thus have existence of strong continuous solutions in the sense of Theorem 6.5 of [Diethelm2010]. These solutions may or may not coincide with solutions of the $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type problem (18). Of course, since the two operators $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}$ and $\operatorname{D}_{\ast 0}^{\alpha}$ coincide almost everywhere in the case of absolutely continuous functions, we have the following statement.
Proposition 3.5.
Consider the unique solution $x\in C[t_{0},t_{0}+h)$ for some $h>0$ of the $\operatorname{D}_{\ast 0}^{\alpha}$-type problem (19). If $x$ is absolutely continuous on $t_{0},t_{0}+h)$, then $x$ satisfies the $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type problem (18) and thus also the Liouville-Weyl problem (12) almost everywhere on $t_{0},t_{0}+h)$.
We remark that, in the case when $f$ is linear in its second argument and autonomous in its first, as the solutions are given by a basis of Mittag-Leffler functions $E_{\alpha}(\lambda(\cdot)^{\alpha})$, the solutions are always absolutely continuous. However, as the following example shows, the assumptions on $f$ being continous in its first argument and Lipschitz in its second, may lead to the $\operatorname{D}_{\ast 0}^{\alpha}$-type problem having solutions which are differentiable nowhere, and thus not absolutely continuous.
Example 3.6.
Consider the Weierstrass function [Samko1994]
$W(t)=\sum_{k=0}^{\infty}\frac{1}{2^{k}}\sin(2^{k}\pi t),\quad t\in[0,1],$
which is continuous but nowhere differentiable on $[0,1]$. This function $W$ is Hölder continuous with exponent $\beta$ for all $\beta\in(0,1)$ and $W(0)=0$. Following the results of [Samko1994], $W$ thus has a continuous Caputo fractional derivative $\operatorname{D}_{\ast 0}^{\alpha}W$ of order $\alpha$ for all $\alpha\in(0,1)$. Then, the function $W$ can be seen as a solution of the initial-value problem
$\operatorname{D}_{\ast 0}^{\alpha}x(t)=f(t),\quad t\in[0,1],\quad x(0)=0$ (20)
with the continuous function $f(t)={}^{\rm C}\!D_{0}^{\alpha}W(t)$. Therefore, the solution of the differential equation (20) is not absolutely continuous. Notice, that $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}W=\operatorname{I}_{0}^{1-\alpha}\operatorname{D}^{1}W$ does not exist, as $\operatorname{D}^{1}W$ also does not exist. The initial value problem (20) can be viewed as a situation in which the solutions of the $\operatorname{D}_{\ast 0}^{\alpha}$-type problem (19) and $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}-$type problem (18) do not coincide.
@@
    The function $x_{0}(t)=\exp(-t)$ does not belong to $X(-\infty,0]$, as it is not bounded on $(-\infty,0]$. Its forcing term does not exist for any $t\geq 0$. Indeed, it holds that
@@
    The function $x_{0}$ is continuous and bounded everywhere and its derivative exists everywhere except at t=1t=-1 and is bounded in the neighborhood of 0. As it is continuously differentiable everywhere but at t=1t=-1, with the singularity at t=1t=-1 being integrable, it is also absolutely continuous.